Vector analysis?A vector A has a magnitude of 53.7 m and points in
Answer #1

A = 53.7 m and 21.5 deg below x-axis

B = 73.9 m and 51.2 deg above x-axis

C = A + B

C = Cx + Cy = Ax + Ay + Bx + By = (Ax + Bx) i + (Ay + By) j

C = (53.7*cos 21.5 deg + 73.9*cos 51.2 deg) i + (-53.7*sin 21.5 deg + 73.9*sin 51.2 deg) j

C = 96.269 i + 37.911 j

|C| = sqrt (96.269^2 + 37.911^2) = 103.464 m = 1.03*10^2 (option B)

Angle = arctan (37.911/96.269) = 21.494 deg = 2.15*10^1 (option D)